### v1

# Geometric Convergence to an Extreme Limit Space with nonnegative scalar curvature

Abstract.
In 2014, Gromov conjectured that sequences of manifolds with nonnegative scalar curvature should have subsequences which converge in some geometric sense to limit spaces with some notion of generalized nonnegative scalar curvature. In recent joint work with Changliang Wang, the authors found a sequence of warped product Riemannian metrics on ${\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$ with nonnegative scalar curvature whose metric tensors converge in the $W^{1,p}$ sense for $p<2$ to an extreme warped product limit space where the warping function hits infinity at two points. Here we study this extreme limit space as a metric space and as an integral current space and prove the sequence converges in the volume preserving intrinsic flat and measured Gromov-Hausdorff sense to this space. This limit space may now be used to test any proposed definitions for generalized nonnegative scalar curvature. One does not need expertise in Geometric Measure Theory or in Intrinsic Flat Convergence to read this paper.
Key words and phrases:
Scalar Curvature, Intrinsic Flat Convergence
Prof. Sormani is partially supported by NSF DMS #1006059.
1. Introduction
Gromov introduced the idea of geometric convergence for a sequence of compact Riemannian manifolds, $(M_{j},g_{j})$, by viewing them as metric spaces, $(M_{j},d_{j})$, where $d_{j}$ is the Riemannian distance, and finding limits that are metric spaces, $(M_{\infty},d_{\infty})$ [13]. He defined the Gromov-Hausdorff (GH) distance between metric spaces, by taking an infimum over all possible distance preserving maps, $\varphi_{j}:M_{j}\to Z$, over all possible compact metric spaces, $Z$, of the Hausdorff distance between the images in $Z$:
(1) $d_{GH}((X_{1},d_{1}),(X_{2},d_{2}))=\inf\,d_{H}^{Z}(\varphi_{1}(X_{1}),\varphi_{2}(X_{2})).$
Note that if one has a sequence of distances, $d_{j}$, all defined on the same compact metric space, $X$, and $d_{j}$ converge uniformly to $d_{\infty}$ as functions, then $(X,d_{j})\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}(X,d_{\infty})$. Gromov-Hausdorff convergence has the advantage of not requiring all the spaces to be the same. Gromov proved a beautiful compactness theorem for GH convergence as well [13].
Gromov-Hausdorff convergence has been applied with great success to study sequences of Riemannian manifolds with nonnegative sectional curvature. The GH-limits of such sequences are Alexandrov spaces which are metric spaces satisfying a generalized notion of nonnegative sectional curvature. See the work of Burago-Gromov-Perelman [6] and others. Gromov-Hausdorff convergence has also been applied combined with measure convergence to study sequences of Riemannian manifolds with nonnegative Ricci curvature. The measured GH-limits of such sequences are $CD$ and $RCD$ spaces which are metric spaces satisfying generalized notion of nonnegative Ricci curvature. See the work of Cheeger-Colding, Lott-Villani, Sturm, and Ambrosio-Gigli-Savare in [2] [8] [18] [24].
In [12] and [11], Gromov vaguely conjectured that a sequence of Riemannian manifolds with nonnegative scalar curvature, $\operatorname{Scalar}\geq 0$, should have a subsequence which converges in some geometric sense to a limit space with some generalized notion of “nonnegative scalar curvature”. He suggested that the intrinsic flat convergence defined by Sormani and Wenger in [23] building upon work of Ambrosio-Kirchheim in [3] might be the appropriate notion. Various conjectures have been formulated in this direction and related examples have been constructed. See the survey by Sormani [21] and papers that cite this survey.
It is not yet known what notion of generalized nonnegative scalar curvature works best for the geometric limits of Riemannian manifolds with nonnegative scalar curvature. In [22], Sormani-Tian-Wang constructed a sequence of warped product manifolds with nonnegative scalar curvature with fibres that stretch to infinite length (see Example 1.1 below). They proved this sequence converges smoothly away from the singular fibres to an extreme warped product space (see Example 1.2 below). In this paper we prove that this sequence converges in a geometric way and describe the limit metric space (see Theorem 1.3 and Theorem 6.1 within). This specific limit space can then be used to test various notions of nonnegative scalar curvature.
The sequence of Riemannian warped product manifolds that we explore in this paper was first introduced Sormani-Tian-Wang in [22]:
Example 1.1.
Consider the sequence of warped product Riemannian manifolds, ${\mathbb{S}}^{2}\times_{g_{j}}{\mathbb{S}}^{1}$, which are diffeomorphic to ${\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$, with Riemannian metrics
(2) $g_{j}=g_{{\mathbb{S}}^{2}}+f_{j}^{2}(r)g_{{\mathbb{S}}^{1}}=dr^{2}+\sin^{2}(r)d\theta^{2}+f_{j}^{2}(r)d\varphi^{2}.$
written using fixed $(r,\theta)$ coordinates on ${\mathbb{S}}^{2}$ and fixed $\varphi$ on the ${\mathbb{S}}^{1}$-fibres where
(3) $f_{j}(r)=\ln\left(\frac{1+a_{j}}{\sin^{2}r+a_{j}}\right)+\beta,$
taking fixed $\beta\geq 2$, positive $a_{j}$ decreasing to 00.
Figure 1. Example 1.2 is an extreme limit warped product space, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},g_{\infty})$, in which the circular fibers, $r^{-1}(0)$ and $r^{-1}(\pi)$, have been stretched infinitely. Since the lengths of any curves in these fibers are infinite, we depict these fibers as circles of discrete points. Away from these two singular fibers, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1}\setminus S,g_{\infty})$ is a smooth Riemannian manifold.
In [22], it was proven that this sequence has $\operatorname{Scalar}_{j}\geq 0$ yet as $j\to\infty$ with diameter and volume bounded above. In addition, closed minimal surfaces in these manifolds cannot have area decreasing to 00 as $j\to\infty$.
In addition Sormani-Tian-Wang proved smooth convergence of the $f_{j}\to f_{\infty}$ on compact sets away from the set
(4) $S=r^{-1}(0)\cup r^{-1}(\pi)$
and defined the following extreme limit space using this limit function.
Example 1.2.
Consider the extreme warped product space, ${\mathbb{S}}^{2}\times_{f_{\infty}}{\mathbb{S}}^{1}$, with metric tensor
(5) $g_{\infty}=g_{{\mathbb{S}}^{2}}+f_{\infty}^{2}(r)g_{{\mathbb{S}}^{1}}=dr^{2}+\sin^{2}(r)d\theta^{2}+f_{\infty}^{2}(r)d\varphi^{2}.$
written using $(r,\theta)$ coordinates on ${\mathbb{S}}^{2}$ and $\varphi$ on the ${\mathbb{S}}^{1}$-fibres with the extreme warping function defined by
(6) $f_{\infty}(r,\theta)=\ln\left(\frac{1}{\sin^{2}r}\right)+\beta=-2\ln\sin r+\beta,$
where $\beta\geq 2$ as in Example 1.1. Note that $g_{\infty}$ is smooth away from the singular set
(7) $S=\{(r,\theta,\varphi)\in{\mathbb{S}}^{2}\times{\mathbb{S}}^{1},\text{ such that }r=0\text{ or }r=\pi\}$
so $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1}\backslash S,g_{\infty})$ is a smooth open Riemannian manifold. See Figure 1.
In addition to proving $g_{j}$ converge to $g_{\infty}$ smoothly away from the singular set, $S$, Sormani-Tian-Wang proved that the metric tensors, $g_{j}$, converge to $g_{\infty}$ in $W^{1,p}({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})$ as $j\to\infty$ for $p\in[1,2)$. They also proved the limit space has nonnegative distributional scalar curvature in the sense defined by D.Lee-LeFloch [14] building upon the work of LeFloch-Mardare [16].
In this paper, we study the geometric notions of convergence of this sequence and the geometric properties of the limit space.
Theorem 1.3.
The sequence of warped product Riemannian manifolds in Example 1.1 viewed as metric spaces, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{j})$, converges in the uniform and Gromov-Hausdorff sense to the metric space $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$ which is the metric completion of the extreme limit space, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1}\setminus S,d_{g_{\infty}})$, of Example 1.2. This limit space is a compact metric space homeomorphic to the isometric product space, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{0})$. In addition the singular set $S$ inside this limit space has Hausdorff dimension $1$ but the singular fibres are unrectifiable, $\mathcal{H}^{1}_{d_{\infty}}(S)=\infty$. Finally ${\mathcal{H}}^{3}_{d_{\infty}}({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})=(2\pi)^{2}(2\beta+4-2\ln 4).$
We also prove volume preserving intrinsic flat convergence of our sequence to a subset of the Gromov-Hausdorff limit whose metric completion is the Gromov-Hausdorff limit. See Theorem 6.1 stated within. These theorems allow others to then test various geometric definitions of generalized nonnegative scalar curvature on this limit space.
Note that this extreme limit space is not an Alexandrov Space with curvature bounded from below, nor is it $CD(3,K)$-space, nor an RCD space since the Ricci curvature of $g_{\infty}$ is not bounded from below on ${\mathbb{S}}^{2}\times{\mathbb{S}}^{1}\backslash S$ as shown in [22]. It is outside the class of $C^{0}$ Riemannian manifolds studied by Bamler and Burkhardt-Guim [4] [7] and by Gromov and Chao Li in [11][17] but still might be studied using their notions of generalized scalar curvature. It has lower regularity than the spaces studied by D.Lee-LeFloch [14] and by M.C.Lee-Topping [15]. It is also outside of the class of Riemannian manifolds with isolated singularities studied by Li-Mantoulidis [17], because the singular set $S$ does not consist of isolated points. We hope it will provide new insight into the full class of possible limit spaces.
This paper is written to be easily read by those who have not studied geometric notions of convergence before. We begin with a background section reviewing the properties of Example 1.1. In Section 3, we prove pointwise convergence of the distance functions, $d_{j}$ to $d_{\infty}$ and prove the claimed homeomorphism. In Section 4 we review GH convergence and prove uniform and GH convergence of $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{j})$ to $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$ applying a theorem of Perales-Sormani from [20]. In Section 5 we prove the metric completion of the smooth limit away from the singular set is $(Sph^{2}\times{\mathbb{S}}^{1},d_{\infty})$ and control the Hausdorff measures. At the end of this section we complete the proof of Theorem 1.3. In Section 6, we state and prove intrinsic flat convergence relying on readers to consult the work of Perales-Sormani for backgroun [20]. In Section 7 we discuss open problems.
2. Background
In this section we review the properties of the sequence of Riemannian warped product manifolds of Example 1.1.
2.1. Warping functions of Example 1.1
The following was proven by Sormani-Tian-Wang in [22]:
Lemma 2.1.
For $f_{j}(r)$ and $g_{j}$ as defined in Example 1.1 we have
(8) $f_{j}(r)\leq f_{j+1}(r)$
and, for all tangent vectors, $V$, we have
(9) $g_{j}(V,V)\leq g_{j+1}(V,V).$
Lemma 2.2.
The sequence $f_{j}$ as defined in Example 1.1 is unbounded in the sense that
(10) $\lim_{j\to\infty}f_{j}(0)\to\infty\textrm{ and }\lim_{j\to\infty}f_{j}(\pi)\to\infty.$
Proposition 2.3.
The sequence $g_{j}$ as defined in Example 1.1 converges smoothly away from $S$ to $g_{\infty}$ of Example 1.2. That is, on compact subsets $K\subset({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\setminus S$ we have
(11) $|f_{j}-f_{\infty}|_{C^{\infty}(K)}\to 0\textrm{ and }|g_{j}-g_{\infty}|_{g_{\infty},C^{\infty}(K)}\to 0$
2.2. Volumes of Example 1.1
The following proposition was proven by Sormani-Tian-Wang in [22]:
Proposition 2.4 (Volumes).
For ${\mathbb{S}}^{2}\times_{f_{j}}{\mathbb{S}}^{1}$ as defined in Example 1.1, the volumes are uniformly bounded from above:
(12) $\operatorname{Vol}_{j}({\mathbb{S}}^{2}\times_{g_{j}}{\mathbb{S}}^{1})\leq 4\pi^{3}\beta.$
2.3. Lengths on Example 1.1
As for any Riemannian manifold, the length of a curve is defined by:
Definition 2.5 (Length of a Curve).
Given a continuously differentiable curve $c:[a,b]\to{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$, we define the length of $c$ as measured by the metric $g$ as
[…]
where $\dot{c}$ denotes the derivative with respect $t$. For the sequence, $g_{j}$, of Example 1.1, we will write, $L_{j}=L_{g_{j}}$.
Sormani-Tian-Wang proved the following lemma in [22]:
Lemma 2.6.
The ${\mathbb{S}}^{1}$ fibres above the poles in Example 1.1 parametrized by the curves,
(14) $c_{0}(t)=(0,0,t)\textrm{ and }c_{\pi}(t)=(\pi,0,t),$
where $t\in[0,2\pi]$ have lengths, $L_{j}(c_{0})=2\pi f_{j}(0)$ and $L_{j}(c_{\pi})=2\pi f_{j}(\pi)$, which diverge to infinity as $j\to\infty$.
2.4. Distances on Example 1.1
On any Riemannian manifold the distance function is defined as follows:
Definition 2.7 (Distance Function).
Given any Riemannian manifold, $(M,g)$, For any $p,\ q\in M$, define
(15) $d_{g}(p,q)=\inf\{L_{g}(c)\},$
where the infimum is taken over continuously differentiable curves, $c:[0,1]\to M$ with $c(0)=p$ and $c(1)=q$ and and where $L_{g}$ is defined as in Definition 2.5.
For Example 1.1, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},g_{j})$ is a smooth Riemannian manifold and we define $d_{j}=d_{g_{j}}$ as in (15) using curves in ${\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$. For Example 1.2, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1}\setminus S,g_{\infty})$ is a smooth Riemannian manifold and we define $d_{g_{\infty}}$ as in (15) using curves that stay within ${\mathbb{S}}^{2}\times{\mathbb{S}}^{1}\setminus S$ avoiding the singular set, $S$.  [^1] We will define $d_{\infty}$ later in Theorem 3.1 as the pointwise limit of the $d_{j}$.
Sormani-Tian-Wang proved the following three results in [22]:
Lemma 2.8 (Distance Functions are Increasing).
The sequence of distance functions, $d_{j}$, as defined in Definition 2.7 satisfies the inequality:
(16) $d_{j}(p,q)\leq d_{j+1}(p,q)\qquad\forall j\in{\mathbb{N}}.$
Lemma 2.9 (Uniform Bound for Distance Functions).
Given any $p_{1}=(r_{1},\theta_{1},\varphi_{1})$ and $p_{2}=(r_{2},\theta_{2},\varphi_{2})$ in ${\mathbb{S}}^{2}\times_{f_{j}}{\mathbb{S}}^{1}$ as defined in Example 1.1,
(17) $d_{j}(p_{1},p_{2})\leq|r_{1}-r_{2}|+\sin(r_{2})\,d_{{\mathbb{S}}^{1}}(\theta_{1},\theta_{2})+f_{j}(r_{2})\,d_{{\mathbb{S}}^{1}}(\varphi_{1},\varphi_{2}).$
Proposition 2.10 (Uniform Bound for Diameter).
For ${\mathbb{S}}^{2}\times_{f_{j}}{\mathbb{S}}^{1}$ as defined in Example 1.1, we have
(18) $\operatorname{Diam}({\mathbb{S}}^{2}\times_{f_{j}}{\mathbb{S}}^{1})\leq D_{0}=(3+2\beta)\pi.$
2.5. The Isometric Product Space
The isometric product of a sphere and a circle is defined as follows:
Example 2.11.
The isometric product tensor on ${\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$ is
(19) $g_{0}=g_{{\mathbb{S}}^{2}}+g_{{\mathbb{S}}^{1}}=dr^{2}+\sin^{2}(r)d\theta^{2}+d\varphi^{2}.$
written using $(r,\theta)$ coordinates on ${\mathbb{S}}^{2}$ and $\varphi$ on the ${\mathbb{S}}^{1}$-fibres. The isometric product distance,
(20) $d_{0}(p,q)=d_{g_{0}}(p,q)=\inf\{L_{g_{0}}(c)\},$
defined as in Definition 2.7.
The following lemma follows easily from Definition 2.5, Definition 2.7, the definition of $g_{0}$ in Example 2.11, and the definition of $g_{j}$ with $f_{j}\geq\beta\geq 1$ in Example 1.1:
Lemma 2.12.
The distances on Example 1.1 and the distances on the isometric product in Example 2.11 satisfy
(21) $d_{j}(p,q)\geq d_{0}(p,q)\quad\forall p,q\in{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}.$
The following two lemmas are well known:
Lemma 2.13.
The isometric product space on $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{0})$ satisfies the Pythagorean Theorem:
(22) $d_{0}((r_{p},\theta_{p},\varphi_{p}),(r_{q},\theta_{q},\varphi_{q}))=\sqrt{d_{{\mathbb{S}}^{2}}((r_{p},\theta_{p}),(r_{q},\theta_{q}))^{2}+d_{{\mathbb{S}}^{1}}(\varphi_{p},\varphi_{q})^{2}}.$
so
(23) $d_{0}((r_{p},\theta_{p},\varphi_{p}),(r_{q},\theta_{q},\varphi_{q}))\geq d_{{\mathbb{S}}^{2}}((r_{p},\theta_{p}),(r_{q},\theta_{q}))\geq|r_{p}-r_{q}|$
and
(24) $d_{0}((r_{p},\theta_{p},\varphi_{p}),(r_{q},\theta_{q},\varphi_{q}))\geq d_{{\mathbb{S}}^{1}}(\varphi_{p},\varphi_{q}).$
Lemma 2.14.
The isometric product space on $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{0})$ is compact.
The following new lemma will be useful to us later:
Lemma 2.15.
Fix $p=(r_{p},\theta_{p},\varphi_{p})\in({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\backslash S$ and choose any
(25) ${\delta\in\left(0,\min\left\{\tfrac{r_{p}}{2},\tfrac{\pi-r_{p}}{2}\right\}\right)}$
For any $q\in{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$ such that $d_{0}(p,q)<\delta$, we have
(26) $d_{{\mathbb{S}}^{1}}(\theta_{p},\theta_{q})<\frac{\delta}{\sin(r_{p}/2)}.$
Proof.
By symmetry, we may assume $r_{p}\in(0,\pi/2]$. Let $c:[0,1]\to({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},g_{{\mathbb{S}}^{2}}+g_{{\mathbb{S}}^{1}})$ be the geodesic connecting $p$ and $q$. In coordinate we have
(27) $c(t)=(r(t),\theta(t),\varphi(t))\text{ for }t\in[0,1].$
Next, we prove that
(28) $r(t)\in(r_{p}/2,3r_{p}/2)\text{ for }t\in[0,1].$
By (23), we have for $t\in[0,1]$
(29) $|r_{p}-r(t)|\leq d_{{\mathbb{S}}^{2}}((r_{p},\theta_{p}),(r(t),\theta(t)))\leq d_{0}(p,c(t))<\delta,$
which by $\delta<r_{p}/2$ and the triangle inequality implies
(30) $r_{p}/2<r_{p}-\delta<r(t)<r_{p}+\delta<3r_{p}/2.$
Moreover, since $r_{p}\in(0,\pi/2]$, we have
(31) $\sin r(t)\geq\sin\left(\frac{r_{p}}{2}\right)\text{ for }t\in[0,1].$
Through direct calculation we have
[…]
which completes the claim. ∎
3. Pointwise Convergence and Homeomorphism
In this section, we introduce the metric space, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$, and prove it is homeomorphic to the isometric product space in Theorem 3.1. We will later prove in Theorem 4.1 that this metric space is the $GH$ limit of $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{j})$. Recall that $d_{j}=d_{g_{j}}$ are the Riemannian distances as in Definition 2.7 of the sequence of Riemannian manifolds in Example 1.1.
Theorem 3.1.
There is a metric space $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$ defined by taking the pointwise limit of the monotone increasing sequence
(33) $d_{\infty}(x,y)=\lim_{j\to\infty}d_{j}(x,y)=\sup_{j\in{\mathbb{N}}}d_{j}(x,y)$
where $d_{j}$ are the distance functions, of the warped product Riemannian manifolds in Example 1.1 as in Definition 2.7. Furthermore this pointwise limit space, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$, is homeomorphic to the isometric product space, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$, of Example 2.11.
Remark 3.2.
We do not yet know how $d_{\infty}$ of Theorem 3.1 is related to $g_{\infty}$ of Example 1.2. We will see this relationship later when we prove Theorem 5.1.
We break the proof of Theorem 3.1 into two propositions that are proven within this section. First we prove pointwise convergence in Proposition 3.3. Then we prove various useful lemmas about $d_{\infty}$ and a couple general technical lemmas. Finally, we prove the homeomorphism in Proposition 3.9 using these lemmas.
3.1. Pointwise Convergence
First we show that the pointwise limit in Theorem 3.1 exists and is a definite metric space:
Proposition 3.3.
The sequence of Riemannian distance functions, $d_{j}=d_{g_{j}}$, of Example 1.1 defined as in Definition 2.7 are monotone increasing and converge pointwise as in (33) to a distance function, $d_{\infty}$, on ${\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$
Proof.
Fix $p,q\in{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$. We have already shown the $d_{j}(p,q)$ is a monotone increasing sequence of real numbers in Lemma 2.8 that is uniformly bounded above by the uniform upper bound, $D_{0}$, on the diameter of these spaces as seen in Proposition 2.10. By the Monotone Convergence Theorem of sequences of real numbers we have (33).
To see that $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$ defines a definite metric space of diameter $\leq D_{0}$, see, for example, Lemma 2.5 of [20] by Perales-Sormani. ∎
3.2. Estimates on Distances
In this section we prove some useful bounds on the pointwise limit, $d_{\infty}$, of Theorem 3.1 that we will apply in a few places later. See Figure 2.
Figure 2. The pointwise limit space, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$, is a metric space in which the distances between points, even those in the circular fibers, $r^{-1}(0)$ and $r^{-1}(\pi)$, can be estimated using Lemma 3.4.
Lemma 3.4.
Given any point $p_{1}=(r_{1},\theta_{1},\varphi_{1})\in{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$ and any point $p_{2}=(r_{2},\theta_{2},\varphi_{2})\in{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}\setminus S$ so that $r_{2}\in(0,\pi)$, we have
(34) $d_{\infty}(p_{1},p_{2})\leq|r_{1}-r_{2}|+\sin(r_{2})\,d_{{\mathbb{S}}^{1}}(\theta_{1},\theta_{2})+f_{\infty}(r_{2})\,d_{{\mathbb{S}}^{1}}(\varphi_{1},\varphi_{2}).$
Proof.
Recall that in [22], Sormani-Tian-Wang estimated $d_{j}(p_{1},p_{2})$ using Definition 2.7 and a specific curve running between the points as depicted in Figure 2. The curve is the concatenation of three curves:
(35) $d_{j}(p_{1},p_{2})\leq L_{g_{j}}(C_{1})+L_{g_{j}}(C_{2})+L_{g_{j}}(C_{3})$
where the curve, $C_{1}$, runs radially from $p_{1}=(r_{1},\theta_{1},\varphi_{1})$ to a point $(r_{2},\theta_{1},\varphi_{1})$ , and then $C_{2}$ runs in the $\theta$ direction to $(r_{2},\theta_{2},\varphi_{1})$ and then $C_{3}$ runs around the fiber to $p_{2}=(r_{2},\theta_{2},\varphi_{2})$. Estimating these lengths using
(36) $g_{j}=dr^{2}+\sin^{2}(r)d\theta^{2}+f_{j}^{2}(r)d\varphi^{2},$
they got
(37) $d_{j}(p_{1},p_{2})\leq|r_{1}-r_{2}|+\sin(r_{2})\,d_{{\mathbb{S}}^{1}}(\theta_{1},\theta_{2})+f_{j}(r_{2})\,d_{{\mathbb{S}}^{1}}(\varphi_{1},\varphi_{2}).$
By Lemma 2.1, we know $f_{j}(r_{2})\leq f_{\infty}(r_{2})$. Substituting this into (37), we have
(38) $d_{j}(p_{1},p_{2})\leq|r_{1}-r_{2}|+\sin(r_{2})\,d_{{\mathbb{S}}^{1}}(\theta_{1},\theta_{2})+f_{\infty}(r_{2})\,d_{{\mathbb{S}}^{1}}(\varphi_{1},\varphi_{2}).$
Our lemma follows by taking the pointwise limit as $j\to\infty$ of this last inequality. ∎
If we have a pair of points which both lie in the singular set, or are close to the singular set, we use the triangle inequality to estimate the distance between them:
Lemma 3.5.
Given any pair of points, $p=(r_{p},\theta_{p},\varphi_{p})$ and $q=(r_{q},\theta_{q},\varphi_{q})$ in ${\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$ and any $r_{0}\in(0,\pi)$, we have
(39) $d_{\infty}(p,q)\leq|r_{p}-r_{0}|+|r_{q}-r_{0}|+\sin(r_{0})\,d_{{\mathbb{S}}^{1}}(\theta_{p},\theta_{q})+f_{\infty}(r_{0})\,d_{{\mathbb{S}}^{1}}(\varphi_{1},\varphi_{2})$
Proof.
Given any pair of points, $p=(r_{p},\theta_{p},\varphi_{p})$ and $q=(r_{q},\theta_{q},\varphi_{q})$ in ${\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$. Let $p^{\prime}=(r_{0},\theta_{p},\varphi_{p})$ and $q^{\prime}=(r_{0},\theta_{q},\varphi_{q})$. By Lemma 3.4 applied three times we have,
(40) $d_{\infty}(p,p^{\prime})$$\leq$ $|r_{p}-r_{0}|+0+0,$
(41) $d_{\infty}(p^{\prime},q^{\prime})$$\leq$ $0+\sin(r_{0})\,d_{{\mathbb{S}}^{1}}(\theta_{p},\theta_{q})+f_{\infty}(r_{0})\,d_{{\mathbb{S}}^{1}}(\varphi_{1},\varphi_{2}),$
(42) $d_{\infty}(q,q^{\prime})$$\leq$ $|r_{q}-r_{0}|+0+0\,\,\,\leq\,\,\,\pi/2.$
By the triangle inequality, $d_{\infty}(p,q)$ is bounded above by the sum of these distances. ∎
Proposition 3.6.
The diameter of the pointwise limit metric space satisfies:
(43) $\operatorname{Diam}_{\infty}({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\leq(3+2\beta)\pi.$
Proof.
This follows from Lemma 3.5 taking $r_{0}=\pi/2$ and using
(44) $|r-\pi/2|\leq\pi/2\qquad\forall r\in[0,\pi],$
and
(45) $f_{\infty}(\pi/2)=-2\ln(\sin(\pi/2))+\beta=\beta,$
and the diameter of $\mathbb{S}^{1}$ is $\pi$. ∎
Lemma 3.7.
The distance on pointwise limit space is bounded below by the distances on the isometric product in Example 2.11:
(46) $d_{\infty}(p,q)\geq d_{0}(p,q)\quad\forall p,q\in{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}.$
Proof.
This follows by taking the pointwise limit of the estimate in Lemma 2.12. ∎
When applying Lemma 3.5, the following lemma will be useful to us:
Lemma 3.8.
For any $m\in\mathbb{N}$, there exists a constant $c_{m}>0$ such that
(47) $f_{\infty}(x)<x^{-1/m}+\beta\quad\forall x\in(0,c_{m}).$
Proof.
By the definition of $f_{\infty}$, we need only prove that there exists $c_{m}>0$ such that
(48) $a(x):=-2\ln(\sin x)x^{\frac{1}{m}}<1\quad\forall x\in(0,c_{m}).$
We take the limit using L’hopital’s rule,
$\lim_{x\to 0^{+}}a(x)$ $=-2\lim_{x\to 0^{+}}\dfrac{\ln(\sin x)}{x^{-\frac{1}{m}}}=-2\lim_{x\to 0^{+}}\dfrac{(1/\sin x)\cos x}{\left(-\tfrac{1}{m}\right)x^{\left(-\frac{1}{m}-1\right)}}$
$=2m\lim_{x\to 0^{+}}\left(\dfrac{x}{\sin x}\right)\cdot\lim_{x\to 0^{+}}\left(x^{\frac{1}{m}}\cos x\right)=0.$
By the definition of limit,
(49) $\forall\epsilon>0\,\,\exists\delta_{\epsilon}>0\textrm{ s.t. }|x-0|<\delta_{\epsilon}\implies|a(x)|<\epsilon.$
Taking $\epsilon=1$, we choose $c=c_{m}=\delta_{\epsilon}$ to complete the proof. ∎
3.3. Homeomorphism
In this section, we prove the homeomorphism part of Theorem 3.1 which we state precisely as follows:
Proposition 3.9.
The identity map $F:(S^{2}\times S_{1},d_{0})\to(S^{2}\times S_{1},d_{\infty})$ is a homeomorphism where $d_{0}$ and $d_{\infty}$ are defined as in Theorem 3.1.
Before we prove Proposition 3.9, we state and prove three lemmas: Lemma 3.10, Lemma 3.11, and Lemma 3.12.
Lemma 3.10.
The function $F$ of Proposition 3.9 is continuous at any point $p\in({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\setminus S$.
Proof.
Fix $p=(r_{p},\theta_{p},\varphi_{p})\notin S$. So $r_{p}\in(0,\pi)$ and
(50) $\bar{d}=\bar{d}_{p}=\min\{r_{p},\pi-r_{p}\}>0$
By Lemma 2.13, the closed ball in the isometric product avoids the singular set,
(51) $\bar{B}_{d_{0}}(p,\bar{d}_{p}/2)\cap S=\emptyset.$
For any $\epsilon>0$, we choose
(52) $\delta=\delta_{\epsilon,p}=\min\{\bar{d}_{p}/2,\epsilon/(1+\sigma_{p}+f_{\infty}(r_{p}))\}>0$
where
(53) $\sigma_{p}=\sin(r_{p})/\sin(r_{p}/2).$
Consider any
(54) $q=(r_{q},\theta_{q},\varphi_{q})\in B_{d_{0}}(p,\delta).$
By the Lemma 3.4, and then by the properties of $d_{0}$ in (23) and (24) and Lemma 2.15, we have
(55) $\begin{split}d_{\infty}(p,q)&\leq|r_{p}-r_{q}|+\sin(r_{p})d_{{\mathbb{S}}^{1}}(\theta_{p},\theta_{q})+f_{\infty}(r_{p})d_{{\mathbb{S}}^{1}}(\varphi_{p},\varphi_{q})\\ &\leq d_{0}(p,q)+\sin(r_{p})\dfrac{\delta}{\sin(r_{p}/2)}+f_{\infty}(r_{p})d_{0}(p,q)\\ &\leq\delta+\sigma_{p}\delta+f_{\infty}(r_{p})\delta\leq\epsilon.\end{split}$
Lemma 3.11.
The function $F$ of Proposition 3.9 is continuous at $p\in S$.
Proof.
Consider any $p=(r_{p},\theta_{p},\phi_{p})\in S$, so $r_{p}=0$ or $r_{p}=\pi$. Recall
(56) $f_{\infty}(r,\theta)=-2\ln(\sin r)+\beta,$
so, without loss of generality, we can assume $r_{p}=0$.
For any $\epsilon>0$, choose
(57) $\delta=\delta_{\epsilon,p}=\min\left\{\tfrac{1}{2},c,\left(\tfrac{\epsilon}{7+\beta}\right)^{3/2}\right\}>0,$
where $c=c_{3}$ is from Lemma 3.8 with $m=3$ which guarantees that
(58) $f_{\infty}(\delta)\leq\delta^{-\frac{1}{3}}+\beta$
Consider any
(59) $q=(r_{q},\theta_{q},\varphi_{q})\in B_{d_{0}}(\delta).$
Since $d_{0}$ is an isometric product metric and $r_{p}=0$,
(60) $r_{q}=|r_{q}-r_{p}|\leq d_{0}(p,q)<\delta$
and
(61) $|\delta-r_{q}|\leq\delta$
By the triangle inequality and then Lemma 3.4 we have,
(62) $\begin{split}d_{\infty}(p,q)&=d_{\infty}((0,\theta_{p},\phi_{p}),(r_{q},\theta_{q},\varphi_{q}))\\ &\leq d_{\infty}((0,\theta_{p},\phi_{p}),(\delta,\theta_{p},\phi_{p}))+d_{\infty}((\delta,\theta_{p},\phi_{p}),(r_{q},\theta_{q},\varphi_{q}))\\ &\leq\delta+|r_{q}-\delta|+\sin^{2}(\delta)d_{{\mathbb{S}}^{1}}(\theta_{p},\theta_{q})+f_{\infty}(\delta)d_{{\mathbb{S}}^{1}}(\varphi_{p},\varphi_{q})\end{split}$
Since $\sin^{2}(\delta)\leq\delta^{2}$ and $d_{{\mathbb{S}}^{1}}(\theta_{1},\theta_{2})<\pi$, we have the following estimate
(63) $|\sin^{2}(\delta)d_{{\mathbb{S}}^{1}}(\theta_{1},\theta_{2})|\leq\pi\delta^{2}$
By (24), we have
(64) $d_{{\mathbb{S}}^{1}}(\varphi_{1},\varphi_{2})\leq d_{0}(p,q)\leq\delta.$
Combining this with (58) we have
(65) $f_{\infty}(\delta)\,d_{{\mathbb{S}}^{1}}(\varphi_{p},\varphi_{q})\leq(\delta^{-1/3}+\beta)\,\delta=\delta^{2/3}+\beta\delta$
Combining these with (61) and then using $\delta<1$, we have
(66) $d_{\infty}(p,q)\leq 2\delta+\delta^{2}\pi+\delta^{2/3}+\beta\delta\,\,\leq(7+\beta)\delta^{2/3}\leq\epsilon.$
Thus $F$ is continuous at singular points, $p\in S$. ∎
Lemma 3.12.
The function $F$ of Proposition 3.9 is bijective and its inverse is Lipschitz one and thus continuous everywhere.
Proof.
The function $F$ is bijective because $F(p)=p$. By Lemma 3.7,
(67) $d_{0}(F^{-1}(p),F^{-1}(q))=d_{0}(p,q)\leq d_{\infty}(p,q).$
So the inverse is Lipschitz $1$ and thus continuous. ∎
We now prove that $F$ is a homeomorphism:
Proof of Proposition 3.9.
This follows immediately from the three lemmas: Lemma 3.10, Lemma 3.11, and Lemma 3.12. ∎
Finally we combine the pointwise convergence of the monotone increasing sequence with the homeomorphism:
Proof of Theorem 3.1.
This follows immediately from Proposition 3.3 and Proposition 3.9. ∎
4. Gromov-Hausdorff Convergence to the Pointed Limit Space
In this section we review Gromov-Hausdorff (GH) convergence and then prove the following GH convergence theorem which will later be applied to prove Theorem 1.3.
Theorem 4.1.
The sequence of warped product Riemannian manifolds in Example 1.1 viewed as metric spaces, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{j})$, converges in the uniform and Gromov-Hausdorff sense to the pointed limit space $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$ which is a compact metric space homeomorphic to the isometric product space, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{0})$.
4.1. Review of Gromov-Hausdorff Convergence
Recall the following definition of the Gromov-Hausdorff distance between compact metric spaces, first introduced by Edwards in [9] and then rediscovered and studied extensively by Gromov in [10].
Definition 4.2.
Given a pair of metric spaces, $(X_{a},d_{a})$ and $(X_{b},d_{b})$, the Gromov-Hausdorff distance between them is
(68) $d_{GH}\left((X_{a},d_{a}),(X_{b},d_{b})\right)=\inf\{d_{H}^{Z}\left(f_{a}(X_{a}),f_{b}(X_{b})\right)\}$
where the infimum is taken over all common metric spaces, $(Z,d_{Z})$, and over all distance preserving maps,
(69) $f_{a}:(X_{a},d_{a})\to(Z,d_{Z})\textrm{ and }f_{b}:(X_{b},d_{b})\to(Z,d_{Z}).$
Here $d_{H}^{Z}$ denotes the Hausdorff distance,
(70) $d_{H}^{Z}\left(A,B\right)=\inf\{R\,:\,A\subset T_{R}(B)\textrm{ and }B\subset T_{R}(A)\}$
between subsets $A,B\subset Z$ which is defined using tubular neighborhoods of a given radius $R$,
(71) $T_{R}(A)=\{z\in Z\,:\,\exists p\in A\,s.t.\,d_{Z}(p,z)<R\}.$
The Gromov-Hausdorff distance is definite in the sense that
(72) $d_{GH}\left((X_{a},d_{a}),(X_{b},d_{b})\right)=0$
iff there is a bijection, $\Psi:X_{a}\to X_{b}$, which is distance preserving:
(73) $d_{b}(\Psi(p),\Psi(q))=d_{a}(p,q)\quad\forall p,q\in X_{a}.$
A sequence of compact metric spaces, $(X_{j},d_{j})$ converges in the Gromov-Hausdorff sense to the compact metric space $(X_{\infty},d_{\infty})$ iff
(74) $d_{GH}\left((X_{j},d_{j}),(X_{\infty},d_{\infty})\right)\to 0.$
For those who would like to learn more about the notion of GH convergence we recommend the book of Burago-Burago-Ivanov [5] and Gromov’s original book [13]. The only result we need for this paper is the following theorem proven by Perales-Sormani in [20].
Theorem 4.3 (Perales-Sormani Monotone to GH).
Given a compact Riemannian manifold, $(M^{m},g_{0})$, possibly with boundary, with a monotone increasing sequence of Riemannian metric tensors $g_{j}$ such that
(75) $g_{j}(V,V)\geq g_{j-1}(V,V)\qquad\forall V\in TM$
with uniform bounded diameter,
(76) $\operatorname{Diam}_{g_{j}}(M)\leq D_{0}.$
Then the induced length distance functions $d_{j}:M\times M\to[0,D_{0}]$ are monotone increasing and converge pointwise to a distance function, $d_{\infty}:M\times M\to[0,D_{0}]$ so that $(M,d_{\infty})$ is a metric space.
If the metric space $(M,d_{\infty})$ is a compact metric space, then $d_{j}\to d_{\infty}$ uniformly and
(77) $(M,d_{j})\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}(M,d_{\infty}).$
This theorem of Perales-Sormani was inspired by early work of Sormani-Tian towards proving GH convergence of our specific monotone sequence. It seemed that many aspects of our planned proof did not use anything more than monotonicity. When Tian chose to leave academia, Sormani decided to prove this more general theorem with Perales and then apply it within this paper.
It is worth noting that without compactness of the limit space, there need not be GH convergence to the pointwise limit space. See, for example, Example 2.6 and Remark 3.8 of [20] by Perales-Sormani.
4.2. Proof of GH Convergence
With our previous results about the pointwise limit, we are now easily able to prove GH convergence:
Proof of Theorem 4.1.
Since the sequence in Example 1.1 is monotone increasing we can apply the Perales-Sormani Monotone to GH Convergence Theorem (stated above as Theorem 4.3). We need only show the pointwise limit space, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$ is compact to obtain uniform and GH convergence of $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{j})$ to $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$.
By our Theorem 3.1, we know the pointwise limit, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$, is homeomorphic to the isometric product metric space $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{0})$. Since ${\mathbb{S}}^{2}$ and ${\mathbb{S}}^{1}$ are compact metric spaces, this isometric product space is compact. Since homeomorphisms preserve compactness, our pointwise limit space, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$, is also compact. ∎
5. The Metric Completion of the Smooth Limit Away from the Singular Set is the GH Limit
In this section, we consider the extreme limit space of Example 1.2 with the singular set removed,
(78) $(({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\setminus S,g_{\infty})$
as a smooth open Riemannian manifold with metric tensor, $g_{\infty}$, that is the smooth limit of $g_{j}$ of Example 1.1 away from the singular set, $S$. It has a Riemannian distance function, $d_{g_{\infty}}$, as in Definition 2.7 defined using curves that avoid the singular set. This defines a metric space,
(79) $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1}\setminus S,d_{g_{\infty}}).$
In contrast, we have the compact metric space, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$, which is defined in Theorem 3.1 and proven in Theorem 4.1 to be the $GH$ limit of $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1}\setminus S,d_{j})$ where $d_{j}=d_{g_{j}}$ as in Definition 2.7).
Here we review the notion of a metric completion and prove the metric completion part of Theorem 1.3 which we state precisely here as Theorem 5.1.
Theorem 5.1.
The metric completion of the extreme limit space, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1}\setminus S,d_{g_{\infty}})$, of Example 1.2 is isometric to the $GH$ limit, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$, of the sequence $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{j})$ of Example 1.1. In addition the singular set $S$ inside this limit space has Hausdorff dimension $1$ and $\mathcal{H}^{1}_{d_{\infty}}(S)=\infty$ and
(80) ${\mathcal{H}}^{3}_{d_{\infty}}({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})=(2\pi)^{2}(2\beta+4-2\ln 4).$
In general the metric completions of smooth limits away from singular sets need not be isometric to the $GH$ limits of the sequences. See, for example, the cinched examples in Allen-Sormani [1].
5.1. Metric Completions
Following Munkres’s pointset topology textbook (cf. [19]) we recall:
Definition 5.2.
Given a metric space, $(X,d)$, the metric completion, $(\bar{X},d)$, is the unique complete metric space, $(Y,d_{Y})$, such that there exists a distance preserving embedding $h:X\to Y$ such that the closure, $\overline{h(X)}$, includes all points in $Y$.
In our setting the space $X=M\setminus S$ and the space $Y=M$ where $M={\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$ and $h:X\to Y$ is the identity map. Our $(Y,d_{Y})$ is the GH and uniform limit space, $(M,d_{\infty})$ which is compact and thus complete. Our $(X,d_{X})$ is $(M/setminusS,d_{g_{\infty}})$. To prove Theorem 5.1, we need only show the identity map
(81) $h:(M/setminusS,d_{g_{\infty}})\to(M,d_{\infty})$
is a distance preserving and that
(82) $\forall p\in M,\exists q_{i}\in M\setminus S\textrm{ such that }d_{\infty}(p,q_{i})\to 0.$
5.2. Lemmas about Monotone Increasing Sequences
Before proving the theorem we prove a pair of general lemmas about monotone increasing metric tensors converging smoothly away from a singular set that can be applied in other settings as well.
Lemma 5.3.
Given any sequence of continuous metric tensors converging pointwise, $g_{j}\to g_{\infty}$ on $N=M\setminus S$, such that
(83) $g_{j}(V,V)\leq g_{j+1}(V,V)\leq g_{\infty}(V,V)\quad\forall V\in TN,$
and given any differentiable curve $\gamma:[0,1]\to N=M\setminus S$ we have the following equality
(84) $\lim_{j\to\infty}L_{g_{j}}(\gamma)=L_{g_{\infty}}(\gamma).$
Proof.
Define the functions, $h_{j}:[0,1]\to[0,\infty)$, to be
(85) $h_{j}(t)=\sqrt{g_{j}(\gamma^{\prime}(t),\gamma^{\prime}(t))}.$
By the hypotheses, each $h_{j}$ is a continuous function satisfying,
(86) $h_{j}(t)\leq h_{j+1}(t)\leq h_{\infty}(t)\qquad\forall t\in[0,1],$
and $h_{j}\to h_{\infty}$ pointwise.
By definition of length $L_{g}$ we have
[…]
By the Monotone Convergence Theorem, we have
[…]
thus
(88) $\lim_{j\to\infty}L_{g_{j}}(\gamma)=L_{g_{\infty}}(\gamma).$
Lemma 5.4.
Given any sequence of continuous metric tensors converging pointwise, $g_{j}\to g_{\infty}$ on a connected $N=M\setminus S$, such that
(89) $\operatorname{Diam}_{g_{j}}(M)\leq D_{0}$
and such that
(90) $g_{j}(V,V)\leq g_{j+1}(V,V)\leq g_{\infty}(V,V)\quad\forall V\in TM,$
Then for any any $p,q\in M\setminus S$, we have
(91) $d_{\infty}(p,q)\leq d_{g_{\infty}}(p,q)$
where
(92) $d_{\infty}(p,q)=\lim_{j\to\infty}d_{j}(p,q)$
which exists because it is a monotone increasing sequence with an upper bound, and where
(93) $d_{g_{\infty}}(p,q)=\inf L_{g_{\infty}}(\hat{c})$
where the infimum is over all continuously differentiable curves $\hat{c}:[0,1]\to M\setminus S$ such that $\hat{c}(0)=p$ and $\hat{c}(1)=q$.
Warning: as seen in the cinched examples of Allen-Sormani in [1], it is possible that
(94) $d_{\infty}(p,q)<d_{g_{\infty}}(p,q)$
if the singular set has been cinched to have shorter lengths.
Proof.
By Definition 2.7, we know that,
(95) $d_{j}(p,q)\leq L_{g_{j}}(\hat{c}).$
By Lemma 5.3, which can be applied to our sequence of metric tensors, $g_{j}$, we have
(96) $L_{g_{j}}(\hat{c})\leq L_{g_{\infty}}(\hat{c}).$
So we have
(97) $d_{j}(p,q)\leq L_{g_{\infty}}(\hat{c}).$
Taking $j\to\infty$, by the pointwise convergence of $d_{j}\to d_{\infty}$, we have
(98) $d_{\infty}(p,q)\leq L_{\infty}(\hat{c}).$
We complete the proof by taking the infimum over $\hat{c}$. ∎
5.3. Avoiding Tubular Neighborhoods of Singular Sets
In this section we show that sufficient controls on the $g_{j}$-lengths of curves that lie in compact sets, $K$, that avoid the singular set, $S$, allow one to estimate $d_{g_{\infty}}$ and ultimately to prove $d_{\infty}=d_{g_{\infty}}$ on $M\setminus S$. We prove two general lemmas and then prove two specific propositions related to our Example 1.1 and its limit Example 1.2.
Lemma 5.5.
Suppose we have a metric space with sequence of increasing distance functions, $(M,d_{j})$ with $d_{j}\leq d_{j+1}$, with a uniform upper bound, $D_{0}$ so that $d_{j}\to d_{\infty}$. Assume in addition that there is a subset $S\subset M$ and there are functions,
(99) $\rho_{j}:M\to[0,D_{0}]$
defined by
(100) $\rho_{j}(p)=\inf\{d_{j}(p,z)\,|\,z\in S\}$
then we have monotone convergence:
(101) $\rho_{j}\leq\rho_{j+1}\to\rho_{\infty}$
where
(102) $\rho_{\infty}(p)=\inf\{d_{\infty}(p,z)\,|\,z\in S\}$
Furthermore the tubular neighborhoods,
(103) $T^{j}_{R}(S)=\rho_{j}^{-1}([0,R))$
satisfy
(104) $T^{j}_{R}(S)\supset T^{j+1}_{R}(S)\supset T^{\infty}_{R}(S)$
and, if $S$ is a compact set, we have for any $j\in{\mathbb{N}}\cup\{\infty\}$,
(105) $\bigcup_{R>0}(M\setminus T^{j}_{R}(S)=M\setminus S.$
Proof.
This is an easy exercise. ∎
In the following proposition we see that our sets are exceptionally simple in our example:
Proposition 5.6.
For our particular sequence of $g_{j}$ in Example 1.1 we have
(106) $\rho_{j}(r,\theta,\phi)=\min\{r,\pi-r\}\quad\forall j\in{\mathbb{N}}\cup\{\infty\}.$
Proof.
This follows for all $j\in{\mathbb{N}}$ because
(107) $S=r^{-1}(0)\cup r^{-1}(\pi)$
and because
(108) $g_{j}(\partial r,\partial r)=1.$
Taking $j\to\infty$ we have $\rho_{j}\to\rho_{\infty}$ so it holds for $j=\infty$ as well. ∎
Lemma 5.7.
Suppose the hypotheses of Lemma 5.4 hold and $\rho_{j}:M\to[0,D_{0}]$ are as in Lemma 5.5. Let $K$ be a compact set in $M$ and let
(109) $d_{j,K}(p,q)=\inf L_{g_{j}}(c)$
where the infimum is over $c:[0,1]\to K$ and $c(0)=p$ and $c(1)=q$. Let
(110) $\lambda_{j}(K)=\max|d_{j}(p,q)-d_{j,K}(p,q)|$
where the sup is over $p,q\in K$. Then for all $p,q\in K$,
(111) $|d_{\infty}(p,q)-d_{\infty,K}(p,q)|\leq\limsup_{j\to\infty}\lambda_{j}(K).$
In particular, for all $p,q\in M\setminus S$,
(112) $|d_{\infty}(p,q)-d_{g_{\infty}}(p,q)|\leq\limsup_{j\to\infty}\lambda_{j}(K)$
for any $K\subset M\setminus S$. So for compact $S$ we have
(113) $d_{\infty}(p,q)=d_{g_{\infty}}(p,q)\qquad\forall p,q\in M\setminus S$
if
(114) $\lim_{R\to 0}\limsup_{j\to\infty}\lambda_{j}(M\setminus T^{\infty}_{R}(S))=0.$
Proof.
Since $g_{j}\to g_{\infty}$ on compact sets, we have
(115) $|d_{j,K}(p,q)-d_{\infty,K}(p,q)|<\delta\quad\forall p,q\in K,\,\,\forall j\geq N^{g,K}_{\delta}$
We already know
(116) $|d_{j}(p,q)-d_{\infty}(p,q)|<\delta\quad\forall p,q\in M,\,\,\forall j\geq N^{d}_{\delta}.$
Thus for all $p,q\in K$,
(117) $|d_{\infty}(p,q)-d_{\infty,K}(p,q)|<2\delta+\lambda_{j}(K)$
for all
(118) $j\geq\max\{N^{g,K}_{\delta},N^{d}_{\delta}\}$
Taking the limsup as $j\to\infty$, and then $\delta\to 0$, we get
(119) $|d_{\infty}(p,q)-d_{\infty,K}(p,q)|\leq\limsup_{j\to\infty}\lambda_{j}(K).$
The claim in (112) follows from the fact that
(120) $d_{\infty}(p,q)\leq d_{g_{\infty}}(p,q)\leq d_{\infty,K}(p,q).$
which follows from Lemma 5.4 and $K\subset M\setminus S$. The final claim follows from the exhaustion by complements of tubular neighborhoods proven in Lemma 5.5. ∎
We now apply the above lemma to Example 1.1.
Proposition 5.8.
For our particular sequence of $g_{j}$ in Example 1.1 we have
(121) $\lambda_{j}(M\setminus T^{j}_{R}(S))=3\pi\sin(R)$
where $\lambda_{j}(K)$ is defined as in (110). In addition,
(122) $\limsup_{j\to\infty}\lambda_{j}(M\setminus T^{j}_{R}(S))=3\pi\sin(R).$
In particular, for all $p,q\in M\setminus S$,
(123) $|d_{\infty}(p,q)-d_{g_{\infty}}(p,q)|\leq 3\pi\sin(R)$
if $r(p),r(q)\in[R,\pi-R]$ which gives us
(124) $d_{\infty}(p,q)=d_{g_{\infty}}(p,q)$
Proof.
Fix $R\in(0,\pi/2)$ and note that the set
(125) $K=M\setminus T^{j}_{R}(S)=r^{-1}[R,\pi-R]$
by Proposition 5.6 and thus is compact and connected. Consider $p,q$ achieving the max in (110) so that
(126) $\lambda_{j}(K)=|d_{j}(p,q)-d_{j,K}(p,q)|.$
Let $c_{j}:[0,1]\to M$ be a $g_{j}$-geodesic with $c_{j}(0)=p$ and $c_{j}(1)=q$ that achieves
(127) $L_{g_{j}}(c)=d_{j}(p,q).$
We claim the geodesic $c_{j}(0,1)$ intersects $r^{-1}(0)\cup r^{-1}(\pi)$ at either an empty set or a single point. Suppose not. Then it intersects at two points. If one of the points is in $r^{-1}(0)$ and the other is in $r^{-1}(\pi)$, then it is no longer minimizing past these points because its projection to the sphere has begun to wrap around and a more efficient geodesic can be chosen. If both lie in $r^{-1}(0)$ or both in $r^{-1}(\pi)$, then the projection of $c_{j}$ to the sphere has a closed loop. We could rotate this loop and its lift in the theta direction by a fixed amount without changing the length of the lift. This would give a new curve that is still continuous but not smooth of minimal length, which is not possible in a smooth Riemannian manifold.
For any $\epsilon>0$, we can find a piecewise smooth curve $c_{j,\epsilon}:[0,1]\to M\setminus S$ with three smooth components with $c_{j,\epsilon}(0)=p$ and $c_{j,\epsilon}(1)=q$ such that
(128) $d_{j}(p,q)\leq L_{g_{j}}(c_{j,\epsilon})=d_{j}(p,q)+\epsilon.$
To do this we need only remove the component net to the single point in $S$ and add a smooth arc with constant $r$ to reconnect them. Since these curves avoid $r^{-1}(0)$ and $r^{-1}(\pi)$, we can now write
(129) $c_{j,\epsilon}(s)=(r_{j,\epsilon}(s),\theta_{j,\epsilon}(s),\phi_{j,\epsilon})$
where each component is a piecewise smooth function. Note that we can guarantee that $\theta_{j,\epsilon}$ is monotone in $s$ on each of the three smooth components of the curve so that
(130) $\int_{0}^{1}|\theta_{j,\epsilon}^{\prime}(s)|/,ds\leq 3\pi.$
Let
(131) $I_{R}=\{s\,:\,r_{j,\epsilon}(s)<R\}\textrm{ and }I^{\pi-R}=\{s\,:\,r_{j,\epsilon}(s)>\pi-R\}$
and define
(132) $r_{R,j,\epsilon}(s)=\left\{\begin{array}[]{ll}R&s\in I_{R}\\ r_{j,\epsilon}(s)&s\in[0,1]\setminus(I_{R}\cup I^{\pi-R})\\ \pi-R&s\in I^{R}\end{array}\right.$
Let
(133) $c_{R,j,\epsilon}(s)=(r_{R,j,\epsilon}(s),\theta_{j,\epsilon}(s),\phi_{j,\epsilon}(s)).$
So $c_{R,j,\epsilon}:[0,1]\to r_{j}^{-1}[R,\pi-R]=K$ and $c_{R,j,\epsilon}(0)=p$ and $c_{R,j,\epsilon}(1)=q$ so
(134) $d_{j,K}(p,q)\leq L_{g_{j}}(c_{R,j,\epsilon}).$
Thus by (126) and $d_{j,K}(p,q)\geq d_{j,K}(p,q)$, we know
(135) $\lambda_{j}(K)$$\leq$ $L_{g_{j}}(c_{R,j,\epsilon})-L_{g_{j}}(c_{j,\epsilon})$
[…]
(137) $\leq$ $\int_{I_{R}\cup I^{\pi-R}}|c_{R,j,\epsilon}^{\prime}(s)|_{g_{j}}-|c_{j,\epsilon}^{\prime}(s)|_{g_{j}}\,\,ds.$
The last line above follows from the fact that these curves agree on $[0,1]\setminus(I_{R}\cup I^{\pi-R})$. By the definition of $g_{j}$ in Example 1.1 we have on $I_{R}\cup I^{R}$
(138) $|c_{R,j,\epsilon}^{\prime}(s)|_{g_{j}}$ $=$ $\sqrt{a_{R}^{2}(s)+b_{R}^{2}(s)+u_{R}^{2}(s)}$
(139) $|c_{j,\epsilon}^{\prime}(s)|_{g_{j}}$ $=$ $\sqrt{a^{2}(s)+b^{2}(s)+u^{2}(s)}$
where
(140) $a_{R}(s)$ $=$ 00 0
(141) $b_{R}(s)$ $=$ $\sin(R)|\theta_{j,\epsilon}^{\prime}(s)|$
(142) $u_{R}(s)$ $=$ $f_{j}(R)|\phi_{j,\epsilon}^{\prime}(s)|$
(143) $a(s)$ $=$ $|r_{j,\epsilon}^{\prime}(s)|$
(144) $b(s)$ $=$ $\sin(r_{j,\epsilon}(s))\,|\theta_{j,\epsilon}^{\prime}(s)|$
(145) $u(s)$ $=$ $f_{j}(r_{j,\epsilon}(s))\,|\phi_{j,\epsilon}^{\prime}(s)|.$
So
(146) $a_{R}(s)\leq a(s)\qquad b_{R}(s)\geq b(s)\qquad u_{R}(s)\leq u(s)$
and thus
(147) $|c_{R,j,\epsilon}^{\prime}(s)|_{g_{j}}$$\leq$ $\sqrt{a^{2}(s)+b_{R}^{2}(s)+u^{2}(s)}$
(148) $|c_{j,\epsilon}^{\prime}(s)|_{g_{j}}$$\geq$ $\sqrt{a^{2}(s)+b^{2}(s)+u^{2}(s)}.$
Taking the difference of these two norms and applying the triangle inequality we get
(149) $|c_{R,j,\epsilon}^{\prime}(s)|_{g_{j}}-|c_{j,\epsilon}^{\prime}(s)|_{g_{j}}\leq|b_{R}(s)-b(s)|\leq|b_{R}(s)|.$
So we can we can estimate the integral using (130) as follows:
[…]
The rest of the claims easily follow. ∎
5.4. Metric Completion Proof
In this section we prove a geeral lemma about metric completions and then a proposition specific to our Example 1.2 and then complete the proof of Theorem 5.1.
Lemma 5.9.
Suppose $(M,d_{\infty})$ is compact and
(152) $d_{g_{\infty}}(p,q)=d_{\infty}(p,q)\quad\forall p,q\in M\setminus S.$
Suppose that for all $z\in S$ there exists $p_{i}\in M\setminus S$ such that $d_{\infty}(p_{i},z)\to 0$. Then $(M,d_{\infty})$ is the metric completion of $(M\setminus S,d_{g_{\infty}})$. In particular, this holds when $S$ is compact.
Proof.
This follows immediately from Definition 5.2. This can be done whenever $S$ is compact because we saw in Lemma 5.5 that
(153) $M\setminus S=\bigcup_{R>0}\,\bigg{(}M\setminus T^{\infty}_{R}(S)\bigg{)}.$
and we can take $R=1/i$. ∎
Proposition 5.10.
In Examples 1.1 and 1.2 for all $z\in S$ we have $p_{i}\in M\setminus S$ such that
(154) $d_{j}(p_{i},z)=1/i\qquad\forall j\in{\mathbb{N}}.$
and so the pointwise limit has
(155) $d_{\infty}(p_{i},z)=1/i\qquad\forall j\in{\mathbb{N}}.$
Proof.
If $z\in S$ then
(156) $z=(0,\theta_{z},\phi_{z})\textrm{ or }z=(\pi,\theta_{z},\phi_{z})$
respectively. We choose
(157) $p_{i}=(1/i,\theta_{z},\phi_{z})\in M\setminus S\textrm{ or }p_{i}=(\pi-1/i,\theta_{z},\phi_{z})\in M\setminus S,$
respectively. By Lemma 3.4, we have
(158) $d_{j}(p_{i},z)\leq 1/i.$
Taking the limit as $j\to\infty$ we have our claim. ∎
5.5. Hausdorff Measure and Dimension of the Singular Set in the Limit Space:
Here we prove the following propositions regarding the Hausdorff measure and dimension of the singular set.
Proposition 5.11.
In $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$ as in Theorem 3.1, the Hausdorff dimension of the singular set $S\subset({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})$ is $1$. In particular, $\mathcal{H}^{3}_{d_{\infty}}(S)=0$. Furthermore $\mathcal{H}^{1}_{d_{\infty}}(S)=\infty$.
Before we can prove this proposition we need the following lemmas.
First we control the distances between points in $r^{-1}(0)$ in the following lemma:
Lemma 5.12.
For any $m\in\mathbb{N}$, we can choose
(159) $\delta_{m}=\min\{c_{m},\tfrac{\pi}{2}\}$
where $c_{m}$ is the constant in Lemma 3.8, with the following property: For any $\varphi_{1},\ \varphi_{2}\in[0,2\pi)$ satisfying
(160) $d_{{\mathbb{S}}^{1}}(\varphi_{1},\varphi_{2})\leq\delta=\delta_{m}$
we have
(161) $d_{\infty}((0,0,\varphi_{1}),(0,0,\varphi_{2}))\leq(3+\beta)\,\delta^{1-\frac{1}{m}}.$
Proof.
Fix $m$ and write $\delta=\delta_{m}$. By Lemma 3.4, we have for all $r\in(0,\pi/2)$
(162) $d_{\infty}((0,0,\varphi_{1}),(0,0,\varphi_{2}))\leq 2r+f_{\infty}(r)\,|\varphi_{1}-\varphi_{2}|\leq 2r+f_{\infty}(r)\,\delta.$
Choosing
(163) $r=d_{{\mathbb{S}}^{1}}(\varphi_{1},\varphi_{2})\leq\delta$
and applying $\delta<\pi/2$, we have
(164) $d_{\infty}((0,0,\varphi_{1}),(0,0,\varphi_{2}))\leq 2\delta+f_{\infty}(\delta)\,\delta.$
Since $\delta=\delta\leq c_{m}$ with Lemma 3.8,
(165) $d_{\infty}((0,0,\varphi_{1}),(0,0,\varphi_{2}))\leq 2\delta+(\delta^{-1/m}+\beta)\,\delta.$
This gives our claim because $\delta<1/2$ implies $\delta^{1/m}<1$ so $\delta\leq\delta^{1-\frac{1}{m}}$ and so
(166) $2\delta+(\delta^{-1/m}+\beta)\,\delta\leq 2\delta^{1-\frac{1}{m}}+\delta^{1-\frac{1}{m}}+\beta\,\delta^{1-\frac{1}{m}}=(3+\beta)\,\delta^{1-\frac{1}{m}}.$
We apply this lemma to prove the following:
Lemma 5.13.
Under the hypotheses of proposition 5.11, the $p$ dimensional Hausdorff measure of $S$ in $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$, denoted $\mathcal{H}^{p}(S)$, is zero for all $p>1$.
Proof.
We need only prove that the $p$ dimensional Hausdorff measure of $S$, denoted $\mathcal{H}^{p}(S)$, is zero for all $p>1$. In fact, we need only show
(167) $\mathcal{H}^{p}(r^{-1}(0))=0\qquad\forall p>1,$
because we can conclude by symmetry that the same holds for $r^{-1}(\pi)$ and because $S=r^{-1}(0)\cap r^{-1}(\pi)$. Recall that
(168) $r^{-1}(0)=\{(0,0,\phi)\,:\,\phi\in{\mathbb{S}}^{1}\}\subset{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}.$
For any $p>1$, there exist a natural number $m=m_{p}>1$ such that
(169) $1+\tfrac{1}{m-1}<p$
Choose $N=N_{m}$ sufficiently large that
(170) $2\pi/N\leq\delta_{m}$
where $\delta_{m}$ is defined as in Lemma 5.12.
Choose a collection of evenly spaced points
(171) $p_{i}=(0,0,2\pi i/N)\in r^{-1}(0).$
By Lemma 5.12,
(172) $r_{N}=d_{\infty}(p_{i},p_{i+1})/2\leq(3+\beta)\left(2\pi N^{-1}\right)^{1-\frac{1}{m}}$
Then the closed $d_{\infty}$-balls of radius $r_{N}$ centered on these points cover $r^{-1}(0)$:
(173) $r^{-1}(0)\subset\bigcup_{j=1}^{N}\bar{B}_{p_{j}}(r_{N}).$
By the definition of Hausdorff measure, there is a constant, $C_{p}$ such that
(174) $\mathcal{H}^{p}(r^{-1}(0))\leq C_{p}\sum_{i=1}^{N}r_{N}^{p}=C_{p}Nr_{N}^{p}.$
By our estimate on $r_{N}$ above,
(175) $\mathcal{H}^{p}(r^{-1}(0))$$\leq$ $C_{p}N(3+\beta)^{p}\left(2\pi N^{-1}\right)^{(1-\tfrac{1}{m})p}$
(176) $\leq$ $C_{p}(3+\beta)^{p}\left(2\pi\right)^{(1-\tfrac{1}{m})p}N^{1-p+p/m}$
We need only show that
(177) $\lim_{N\to\infty}N^{1-p+\tfrac{p}{m}}=0.$
So we need only show
(178) $0>1-p+\tfrac{p}{m}.$
Since our choice of $m=m_{p}$ at the top in (169), is positive, we need only show
(179) $0>m-mp+p$
which follows from $mp-p>m$ and $p(m-1)>m$. Since our choice of $m=m_{p}>1$, we need only show
(180) $p>m/(m-1)$
which follows from our choice of $m=m_{p}$ in (169). Hence $\mathcal{H}^{p}(S)=0$ for all $p>1$. ∎
Lemma 5.14.
Under the hypotheses of Proposition 5.11, the singular set, $S$, has Hausdorff dimension one in $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$. However
(181) $\mathcal{H}^{1}_{d_{\infty}}(r^{-1}(0))=\infty\textrm{ and }\mathcal{H}^{1}_{d_{\infty}}(r^{-1}(\pi))=\infty.$
In fact the continuous curves, $c_{r}:{\mathbb{S}}^{1}\to r^{-1}(0)$, defined by $c_{r}(t)=(r,0,t)$ for $r=0$ and for $r=\pi$ are unrectifiable bijections.
Proof.
We know that for $r=0$ and $r=\pi$ that $c_{r}$ is a bijective homomorphism, by Proposition 3.9. By Lemma 2.6, we know
(182) $L_{g_{j}}(c_{r})=2\pi f_{j}(0)$
which means $c_{r}$ is $d_{j}$-rectifiable and
(183) $2\pi f_{j}(0)=\sup\sum_{i=1}^{N}d_{j}(c_{r}(t_{i-1}),c_{r}(t_{i}))$
where the supremum is over all $N\in{\mathbb{N}}$ and over all partitions,
(184) $P=\{t_{0},t_{1},...,t_{N}\}$
such that
(185) $0=t_{0}<t_{1}<\cdots<t_{N}=2\pi.$
Thus for each $j\in{\mathbb{N}}$, there is $N_{j}\in{\mathbb{N}}$, and a partition,
(186) $P_{j}=\{t_{j,0},t_{j,1},...,t_{j,N}\}$
such that
(187) $\pi f_{j}(0)<\sum_{i=1}^{N}d_{j}(c_{r}(t_{j,i-1}),c_{r}(t_{j,i})).$
Since $d_{\infty}(p,q)\geq d_{j}(p,q)$ for all $p,q\in{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$ we have
(188) $\pi f_{j}(0)<\sum_{i=1}^{N}d_{\infty}(c_{r}(t_{j,i-1}),c_{r}(t_{j,i})).$
We claim $c_{r}$ is not $d_{\infty}$-rectifiable. Assume on the contrary that $c_{r}$ has finite $d_{\infty}$-rectifiable length, $L_{d_{\infty}}(c_{r})$. Then it is larger than the sum and we have
(189) $\pi f_{j}(0)<L_{d_{\infty}}(c_{r}).$
However $\lim_{j\to\infty}f_{j}(0)=\infty$, so this is a contradiction.
Since $c_{0}$ and $c_{\pi}$ are $d_{\infty}$-unrectifiable bijections their images have unbounded Hausdorff measure of dimension one as claimed in (181). Thus ${\mathcal{H}}^{1}(S)=\infty$. Combining this with Lemma 5.13, we see that the Hausdorff dimension of the singular set is $1$. ∎
We can now prove Proposition 5.11:
Proof.
This follows immediately from Lemma 5.13 and Lemma 5.14. ∎
5.6. Volume of the Limit Space
We can now find the volume of the limit space:
Proposition 5.15.
The metric space $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$ as in Theorem 3.1 has volume,
(190) ${\mathcal{H}}^{3}_{d_{\infty}}({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})=(2\pi)^{2}(2\beta+4-2\ln 4).$
Proof.
Since ${\mathcal{H}}^{3}_{d_{\infty}}(S)=0$ we know
(191) ${\mathcal{H}}^{3}_{d_{\infty}}({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})={\mathcal{H}}^{3}_{d_{\infty}}(({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\setminus S,d_{\infty}).$
By Proposition 5.8 we have
(192) $d_{\infty}(p,q)=d_{g_{\infty}}(p,q)\quad\forall p,q\in({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\setminus S.$
so
(193) ${\mathcal{H}}^{3}_{d_{\infty}}(({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\setminus S,d_{\infty})=\operatorname{Vol}_{g_{\infty}}(({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\setminus S).$
By Lemma 3.3 in [22] by Sormani-Tian-Wang which is proven by integration, we have
(194) $\operatorname{Vol}_{g_{\infty}}(({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\setminus S)=(2\pi)^{2}(2\beta+4-2\ln 4).$
5.7. Proof of our Main Theorem
We can now prove Theorem 5.1 and then apply it to complete the proof of Theorem 1.3
Proof of Theorem 5.1.
In Theorem 3.1 we saw that $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$ is homeomorphic to the isometric product, ${\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$, and is thus compact. In Proposition 5.8 we saw that
(195) $d_{\infty}(p,q)=d_{g_{\infty}}(p,q)\quad\forall p,q\in({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\setminus S.$
In Proposition 5.10 we saw that all $p\in{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$ lie in the closure of $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\setminus S$. Thus by Definition 5.2 $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$ is the metric completion of $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})\setminus S,d_{g_{\infty}})$. The two claims regarding Hausdorff measures follow from Proposition 5.15 and Proposition 5.11. ∎
We can now prove Theorem 1.3 by combining our three theorems:
Proof of Theorem 1.3.
This is an immediate consequence of Theorem 3.1, Theorem 4.1, and Theorem 5.1 proven above. ∎
6. Intrinsic Flat Convergence
In this section, we prove the volume preserving intrinsic flat ($\mathcal{VF}$) convergence of our sequence. Intrinsic flat convergence of oriented Riemannian manifolds and more general integral current spaces was introduced by by Sormani-Wenger in [23] using the deep theory of integral currents and Lipschitz tuples on metric spaces developed by Ambrosio-Kirchheim in [3].
Theorem 6.1.
If we view the sequence of Riemannian manifolds $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},g_{j})$ of Example 1.1 as integral current spaces,
(196) $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{j},[[{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}]]),$
then they converge in the volume preserving intrinsic flat sense to $(M_{\infty},d_{\infty},T_{\infty})$ where $d_{\infty}$ is the distance on the Gromov-Hausdorff limit, $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$, and where
(197) $M_{\infty}={\rm{set}}_{d_{\infty}}(T_{\infty})\subset{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}\textrm{ with closure }\overline{M}_{\infty}={\mathbb{S}}^{2}\times{\mathbb{S}}^{1}.$
and where $T_{\infty}$ is an integral current on $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{\infty})$ such that $T_{\infty}=[[{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}]]$ viewed as an integral current on $(M,d_{j})$ which means for any $j\in{\mathbb{N}}$:
[…]
for any tuple of Lipschitz functions,
(199) $\pi_{i}:({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{j})\to\mathbb{R}$
where $\pi_{0}$ is bounded.
We begin with a review of $\mathcal{VF}$ convergence and integral currents and then prove key lemmas before completing the proof of this theorem.
6.1. A Theorem about Intrinsic Flat Convergence
Due to limited space, we have no room to review the Geometric Measure Theory needed to truly understand this part of the paper.
For those who would like to learn enough about the notion of volume preserving intrinsic flat ($\mathcal{VF}$) convergence and integral current spaces to understand the statements of our Theorem 6.1 and the theorem below, we recommend reading the background in the paper of Perales-Sormani [20]. The only result about volume preserving intrinsic flat convergence that we need for this paper is the following theorem proven by Perales-Sormani in [20]:
Theorem 6.2 (Perales-Sormani-Monotone-Convergence).
Given a compact Riemannian manifold, $(M^{m},g_{0})$, possibly with boundary, with a monotone increasing sequence of Riemannian metric tensors $g_{j}$ such that
(200) $g_{j}(V,V)\geq g_{j-1}(V,V)\qquad\forall V\in TM$
with uniform bounded diameter,
(201) $\operatorname{Diam}_{g_{j}}(M)\leq D_{0}.$
Then the induced length distance functions $d_{j}:M\times M\to[0,D_{0}]$ are monotone increasing and converge pointwise to a distance function, $d_{\infty}:M\times M\to[0,D_{0}]$ so that $(M,d_{\infty})$ is a metric space.
If the metric space $(M,d_{\infty})$ is a compact metric space, then $d_{j}\to d_{\infty}$ uniformly and
(202) $(M,d_{j})\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}(M,d_{\infty}).$
If, in addition, $M$ is an oriented Riemannian manifold with uniformly bounded total volume,
(203) $\operatorname{Vol}_{j}(M)\leq V_{0}\textrm{ and }\operatorname{Vol}_{j}(\partial M)\leq A_{0}$
then we have volume preserving intrinsic flat convergence
(204) $(M,d_{j},[[M]])\stackrel{{\scriptstyle\mathcal{VF}}}{{\longrightarrow}}(M_{\infty},d_{\infty},T_{\infty}).$
where “$T_{\infty}=[[{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}]]$ viewed as an integral current structure on $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{j})$” and $M_{\infty}$ satisfies
(205) $M_{\infty}={\rm{set}}_{d_{\infty}}(T_{\infty})\subset M\textrm{ with closure }\overline{M}_{\infty}=M.$
This means
[…]
for any tuple of Lipschitz functions
(207) $\pi_{i}:({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{j})\to{\mathbb{R}}$
with $\pi_{0}$ bounded.
Note that we have already applied the top part of this theorem, which we called the Perales-Sormani Monotone to GH Theorem (stated as Theorem 4.3 above) to prove GH convergence of our sequence. We avoided the full statement then because intrinsic flat convergence part is somewhat complicated, and we wanted to ensure the first part of the paper was easy to ready for those only interested only in GH limits.
This section is more difficult but can be followed easily after reading the background section of [20] by Perales-Sormani. One might even read that entire paper first to truly understand what is proven here. If one accepts Perales-Sormani-Monotone-Convergence Theorem as a black box, then you can still read the proof below.
6.2. Proof of $\mathcal{VF}$ Convergence to the Extreme Limit Space
In this section we will prove the intrinsic flat convergence.
Proof of Theorem 6.1.
We have already proven every hypothesis in the first paragraph of the Perales-Sormani Monotone Convergence Theorem (stated as Theorem 6.2 above) to prove our Gromov-Hausdorff convergence in Theorem 4.1. Sormani-Tian-Wang proved a uniform upper bound on $\operatorname{Vol}_{g_{j}}({\mathbb{S}}^{2}\times{\mathbb{S}}^{1})$ in [22] (which we reviewed in Proposition 2.4). Our manifold, ${\mathbb{S}}^{2}\times{\mathbb{S}}^{1}$, has no boundary. Thus, we have all the hypotheses of Theorem 6.2. So we conclude volume preserving intrinsic flat convergence to
(208) $(M_{\infty},d_{\infty},T_{\infty})$
where $M_{\infty}$ satisfies (205) and “$T_{\infty}=[[{\mathbb{S}}^{2}\times{\mathbb{S}}^{1}]]$ viewed as an integral current structure on $({\mathbb{S}}^{2}\times{\mathbb{S}}^{1},d_{j})$”. ∎
7. Testing Notions of Nonnegative Scalar Curvature
Now that we have a strong understanding of the metric properties of our extreme limit space, this space can be used to test various proposed notions of generalized nonnegative scalar curvature. Here we suggest two particular notions to check, but others may be checked on this extreme example as well.
Remark 7.1.
It would be interesting to explore whether our extreme limit space of Example 1.2 has nonnegative scalar curvature in the sense of prism rigidity as studied by Gromov and Chao Li in [11] and [17]. Since it is known that the extreme limit space is not flat anywhere, one would need only show there are no prisms satisfying the hypotheses of Gromov’s prism property. Note that it has already been shown there are no such prisms in the sequence of Example 1.1 because they have positive scalar curvature, so this might be helpful towards completing a proof. The fact that we have $C^{0}$ convergence on some regions should also be helpful.
Remark 7.2.
It would be interessting to see if one could define a Ricci flow starting from our extreme limit space of Example 1.2 and see if it flows into smooth spaces of positive scalar curvature. One way to do this might be to find the Ricci flow of the sequence of Riemannian manifolds in Example 1.1 which definitely flow into smooth manifolds of positive scalar curvature. Then try to take a limit of these flows and show they converge smoothly for $t>0$ even though they only converge in the GH or SWIF or uniform sense as $t\to 0$. If this works, then our extreme limit space has generalized nonnegative scalar curvature in the sense defined by Bamler and Burkhardt-Guim in [4] and [7]